Mock mid term winter 2019 ences Mailings Review View Help Gramm arly ρ Tell me what you want to do Paragraph Styles Multiple Choice 3) A package falls off a truck that is moving at 30 m/s. Neglecting air resistance, the horizontal speed of the package just before it hits the ground is a. zero b. less than 30 m/s, but larger than zero c. about 30 m/s d. more information needed for an estimate. 4) A bullet is dropped into a river from a very high bridge. At the same time, another bullet is fired from a gun, straight down towards the water. Neglecting air resistance, the acceleration just before striking the water a. is greater for the dropped bullet b. is greater for the fired bullet. c. is the same for each bullet d. depends on how high they started S) It takes 6s for a stone to fall to the bottom of a mine shaft. How deep is the shaft? a. about 60 m b. about 120m c. about 180 m
Answer #1

3) When the packet drops off a moving truck,in vertical direction it starts accelerating due Gravitational attraction, which is an external force.

On the other hand,the package had a horizontal sped of 30 m/s and will continue with that speed during its further motion as predicted by Newton's Second Law of Motion. Hence it's velocity will be 30 m/s ( Answer C)

4) When a bullet falls from bridge it's initial vertical speed = Vo =0m/sec

When bullet is fired from bridge it's initial velocity = Vo m/s ( Vo > 0m/s)

In view of above the bullet which leaves the gun has a final velocity = Vf = Vo + gt which is greater than the bullet in free fall.

5) Time of vertical flight = 6 secs

Distance fallen in 6 secs = 0.5 X g X t^2 = 0.5 x 9.8 m/s^2 x 6^2 = 176.4 m

Thus the answer is C ( about 180 m)